OpenStudy (anonymous):

2 divided by 3+ the square root 3

OpenStudy (anonymous):

\[2/(3+\sqrt{3}) \] Is this the form that you mean? If so then you need to simplify using the conjugate of the denominator. The conjugate is means that I will change the sign between the terms. So the conjugate here would be\[(3-\sqrt{3})\] Now in order to use it, multiply by \[(3-\sqrt{3})/(3-\sqrt{3})\] Then simplify using distribution.